the instructor of an introductory computer course wants to make a wall chart of a computer desktop for her students. to make it as realistic as possible, she would like the desktop icons to be proportional to the size of her 6ft by 6ft chart. if the icons on her 15 inch by 15 inch pc screen are 3/8 inch by 3/8 inch, how large should they be on her wall chart.
Answer by bucky(2189)

This problem uses the process of scaling.
.
The question to ask yourself is how much do I have to enlarge a 15 inch dimension to grow
it to 6 feet. Note that 6 feet is 6*12 = 72 inches. So the question becomes by what factor
do I have to multiply 15 inches to zoom it out to 72 inches. In equation form this becomes:
.
15 * x = 72
.
In which x is the zoom (or magnification or scaling) factor. Solve for x by dividing
both sides of the equation by 15 to get:
.
x = 72/15 = 4.8
.
If you multiply the 15 inch dimension of the computer screen by 4.8 times it will become
the size of the 6 ft wall chart.
.
But if you are multiplying the 15 inch dimension by 4.8 times, you will also have to
multiply the 3/8 inch icon by the same factor when you put it on the wall chart.
.
Note that 3/8 inch is equivalent to 0.375 inches. (You get this by dividing the denominator
8 into the numerator 3.) So the dimension of the icon on the wall chart will be 0.375 inches
times the magnification factor 4.8. The product of 0.375 times 4.8 is 1.8 inches.
Therefore, the size of the icon on the wall chart will be 1.8 inches by 1.8 inches.
.
Hope this helps you to understand the problem, and gives you some insight into the process
of scaling.
.